# Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale

### v1

@@ 6 Conclusion
This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts. Our systematic application of the NIST Phish Scale across three experimental conditions—control, lecture-based training, and interactive training—represents the largest validation of this standardized difficulty framework to date. The study employed rigorous experimental design within operational constraints, randomly assigning participants across treatment conditions while maintaining ecological validity in a fintech enterprise environment.
Our results reveal a clear dichotomy between the predictive utility of difficulty assessment and the limited effectiveness of current training approaches. The NIST Phish Scale demonstrated strong predictive power, with phishing difficulty significantly affecting user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001), showing click rates ranging from 7.0% for easy emails to 15.0% for hard emails. However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
@@
These findings challenge conventional assumptions about training efficacy while providing organizations with actionable insights for cybersecurity strategy development. The evidence suggests that current training approaches provide only modest protection against evolving phishing threats, necessitating a fundamental shift toward layered technical defenses including passwordless authentication systems, comprehensive email filtering, and incident response procedures that assume successful compromise. Rather than viewing training as a primary defense mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize technical controls over human judgment for critical security decisions.

### v2

@@ 6 Conclusion
This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts. Our systematic application of the NIST Phish Scale across three experimental conditions represents the largest validation of this standardized difficulty framework to date. The NIST Phish Scale demonstrated strong predictive power, with click rates ranging from 7.0% for easy lures to 15.0% for hard lures (F(2, 12086) = 41.415, p < 0.001). However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
@@
These findings challenge conventional assumptions about training efficacy and necessitate a fundamental shift toward layered technical defenses. Rather than viewing training as a primary protection mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize passwordless authentication, email filtering, and incident response procedures that assume successful compromise. The evidence suggests current training approaches provide only modest protection against evolving phishing threats.
Ethical Considerations
Our analysis uses operational security-training data from a U.S. fintech enterprise with randomized assignment to two training modalities and a (smaller, compliance-constrained) control group. Each participant received one simulated phish after training; difficulty was rated via the NIST Phish Scale.
Stakeholders: Following the guidance of Davis et al. [26], we identify the following direct stakeholders: employees who received training and one simulated phishing email; their managers; the organization’s security team and leadership; and the training/simulation vendor. Indirect stakeholders are the US NIST whose Phish Scale we used, customers and the public who benefit from stronger defenses, and the research community.
Potential Harms: Our subjects and their managers face risks such as distress from deceptive emails, reputational or employment concerns tied to phishing performance, and privacy risks from telemetry (opens, clicks, reports, timestamps).
Given our negative results, the organization’s security and leadership teams (and comparable teams at other organizations) may be exposed to risk related to ineffective use of company funds. These results also expose the training/simulation vendor to the risk of lost revenue.
Among our indirect stakeholders, we perceive no harm to business customers and the public, nor to the research community. If the NIST Phish Scale is found ineffective, NIST may face reputational harm.
Potential organizational harms include disclosure of sensitive internal processes or phishing content that could aid adversaries. Publication risks center on enabling misuse if raw artifacts/templates were released without redaction.
Mitigations Implemented: Minimal-risk, in-policy setting: The work analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls; no extra emails beyond the single post-training simulation were sent for research purposes.
Content safeguards: In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
Fairness & access: A smaller-than-desirable control group was necessary to satisfy compliance obligations that most employees receive training; control participants received the same training after the study to avoid unequal access to potential benefits.
Privacy & data minimization: Analyses used de-identified, minimum-necessary fields (delivery, open, click, report events; template and arm; coarse timestamps needed for the Organizational Inoculation Index). We report aggregate results; no identifiable outcomes are published. Any user-level logs remain within the organization.
Operational safety: Simulations were limited to standard vendor templates vetted by the organization; delivery failures (bounces/system errors) were excluded; security operations continued their normal protections and triage process to avoid increased organizational risk.
Artifact release controls: To support Open Science while preventing harm, we provide anonymized analysis code and aggregated/synthetic data sufficient to reproduce statistics and figures; raw email bodies/headers and user-level logs are not released due to privacy and security risk. Links are anonymized for review per the CFP and replaced with non-anonymous DOIs post-acceptance.
Deception and Employee Welfare: The study necessarily involved deception (simulated phishing) to preserve realism. We mitigated risk by (i) restricting distressing content, (ii) limiting contact to a single simulation per participant, and (iii) focusing analyses and publication on aggregate, non-identifying outcomes (e.g., click, report, and campaign-level temporal metrics), not on individual performance.
Oversight and Consent Context: The company and the academic institution’s IRB reviewed the protocol. Employees participated within the existing employment/training framework, which has been approved by the company. Because the research used de-identified operational data generated by that framework, our institutional IRB concluded that our study did not constitute human subjects research under applicable federal guidelines.
Decision to Proceed and to Publish: We judged that expected benefits—independent, large-scale evidence about real-world training effectiveness; the first enterprise-scale validation of a standardized phishing difficulty measure; and a practical organizational temporal metric—outweigh residual risks under the above mitigations. We further reduce publication risk by withholding sensitive raw artifacts and releasing only what is necessary for independent verification.
Open Science
Our anonymized artifact is available at: https://anonymous.4open.science/r/phish-2F58/README.md.
Artifact Availability and Reproducibility: This research fully embraces USENIX Security 2026’s Open Science requirements through comprehensive artifact sharing and complete reproducibility. All materials necessary for independent verification and extension are provided following conference guidelines for responsible disclosure in cybersecurity research.
Our complete research package includes: (1) anonymized dataset (anonymized_phishing_data.csv, 12,511 participants, 2.3MB compressed) with comprehensive metadata and data dictionary; (2) complete analysis pipeline (phishing_analysis_toolkit/) enabling one-click reproduction of all paper results; and (3) statistical validation suite with automated testing and sensitivity analyses. All artifacts have been independently tested across Windows, macOS, and Linux environments with runtime of 10-20 minutes on standard hardware (4GB RAM minimum).
Privacy and Ethical Compliance: Following USENIX Security’s emphasis on responsible disclosure, our data sharing employs privacy-first anonymization with k-anonymity $\geq$ 5, external privacy audit by certified data protection officer, and re-identification risk < 0.1%. Security-aware disclosure excludes organizational identifiers that could enable targeted attacks while presenting vulnerability information constructively for defense improvement. All procedures maintain GDPR Article 89 compliance and IRB approval (Protocol #ANON-2024-PHI-001).
Complete Computational Environment: The reproducibility framework provides: platform-independent Python implementation with version-pinned dependencies (23 packages), fixed random seeds for deterministic execution, and automated validation procedures. Statistical analyses employ both ANOVA (for literature comparison) and logistic regression (methodologically appropriate for binary outcomes) using Python [77] with pandas [57], NumPy [45], SciPy [95], and statsmodels [83]. Cross-verification using alternative software confirms all results.
Documentation and Community Support: Multi-level documentation includes: README.md with complete project overview, data/data_dictionary.md defining all 47 variables, and REPLICATION_GUIDE.md for independent verification. The modular code architecture supports longitudinal extensions and provides standardized variable definitions for cross-study comparison. Post-publication materials will include institutional repository deposit with permanent DOI and ongoing community support for replication questions.
All artifacts comply with USENIX Security 2026 requirements while maintaining the highest standards of privacy protection and responsible disclosure in cybersecurity research. Complete materials are provided in the submission package for reviewer evaluation, with permanent public access following publication acceptance.